OpenStudy (anonymous):

Q: The height and the base of a tiangle each measure 2 2/3 (2/3 is expoent ) units. What is the area of the triangle?

OpenStudy (anonymous):

So \[(2^{\frac{ 2 }{ 3 }}\times2^{\frac{ 2 }{ 3 }}) \div 2\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

if \[\frac{ 2 }{ 3 }\] is a exponent

OpenStudy (anonymous):

1.25992105

OpenStudy (anonymous):

what how u get that?

OpenStudy (anonymous):

\[(2^{\frac{ 2 }{ 3 }}\times2^{\frac{ 2 }{ 3 }}) = 2.5198421 \div 2\]

OpenStudy (anonymous):

OK u sure that is the right answer?

OpenStudy (anonymous):

Yeah, I'm using a calculator.

OpenStudy (anonymous):

but is looks small for a triangle

OpenStudy (anonymous):

its less than 2

OpenStudy (anonymous):

About 1.26

OpenStudy (anonymous):

sry it was 3/2

OpenStudy (anonymous):

The exponent?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Hold on...\[(2^{\frac{ 3 }{ 2 }}\times2^{\frac{ 3 }{ 2 }})\div2\]

OpenStudy (anonymous):

4

OpenStudy (anonymous):

answer is 4?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u show the steps